OpenStudy (anonymous):

The length of the hypotenuse of a 30°-60°-90° triangle is 11. What is the perimeter?

OpenStudy (_michaeld22):

\[\frac{ 33 }{ 2 } + \frac{ 11 }{ 2 }\sqrt{3}\]

OpenStudy (anonymous):

very good, it's 18+63−−√

OpenStudy (anonymous):

\[18+6\sqrt{3}\]